OpenStudy (anonymous):

What is the period amplitude and phase shift of one cycle of h(x) = 3/2 sin (2x - pi)

OpenStudy (anonymous):

h(x) = a sin (bx - c) + d Amplitude = |a| = 3/2 Period = 2π/|b| = π Phase Shift = -c/b = π/2 Vertical Shift = d = 0

OpenStudy (anonymous):

thanks!